Two vectors are given by vector a =5.3î+9.5ĵ and vector b =4.4î+2.4ĵ. Find (a)| vector a × vector b |, (b) vector a ⋅ vector b, (c)(vector a + vector b) . vector b , and (d) the component of vector a along the direction of vector b?
Answer #1

axb=| ax ay az |=1(9.5.0-0.2.4)-วั53.0-o-4.4)-k(5.3.2.4-9.5.4.4) = 5.3 9.5 0 4.4 2.4 0 -i (0-0)-j(0-0)-k (12.72-41.8) = {0; 0;-29.08 ; -be ax . brt au . Њут az , b,-5.3 . 4.4 + 9.5 . 2.4 + 0 . 0-23.32+ 22.8 + 0-1153 25 ab b- abr bz+ aby by abz b-9.7 44 11.9 24+0 0 42.68+28.56+0 25 b.-M5 + bú-b2 04.4)2 + (2.4)2 19.36-5.76-0-125 12-2 157 ~ 5.01 1985634456667 0-

as component of a along b = a.b / |b|

a long b = 1153 / 25 / 5.011 = 9.2